OpenStudy (anonymous):

can someone help me solve this equation: 2^2x-2^x=8

OpenStudy (anonymous):

actually my answer is wrong. solve \[u^2-u-8=0\]

OpenStudy (anonymous):

Let X = 2^x: x^2 - x - 8 = 0 X = (1 +/- V33) /2 2^x = (1 +/- V33) /2 Take log2 both sides: => x = log2 (1 + V33) /2

OpenStudy (anonymous):

= 1.75

OpenStudy (radar):

\[2^{3.5}-2^{1.75}=7.9502\] Close but is it close enough??

OpenStudy (radar):

I think so, but I'm retired guvemint

OpenStudy (radar):

Hi robtobey

OpenStudy (anonymous):

No real root. Refer to the attachment.

OpenStudy (anonymous):

Hi radar.

OpenStudy (radar):

That is what Wolframalpha say also, a complex solution. Have a nice day robtobey